# Subtype-Aware Registration of Longitudinal Electronic Health Records

### v1

@@ 3 Simulations
In summary, our algorithm effectively recovered the temporal positions from the simulated data across all six scenarios. The evaluation metrics show similar values regardless of the complexity of the scenario, indicating the robustness of our algorithm. The total runtime per execution of the algorithm on the simulated dataset does not exceed 2 minutes, demonstrating its low computational complexity and potential for practical application on large-scale EHR datasets.
Acute Kidney Injury (AKI) is a common clinical condition characterized by an abrupt decline in kidney function, resulting from various causes such as decreased renal perfusion, direct nephron injury, or urinary tract obstruction [36]. AKI can develop rapidly within hours or days and is associated with significant morbidity and high mortality rates [37]. This syndrome is highly heterogeneous, and the complexity of its pathophysiology poses significant challenges in developing effective patient treatment plans [37].
In this study, we utilize longitudinal EHR data from AKI patients to demonstrate that our proposed method effectively aligns observed EHR records among different subjects who belong to the same cluster. Specifically, we employ unsupervised learning techniques, complemented by clinical evaluations and supervised learning, to validate the effectiveness of our subtype-aware EHR registration method.
4.1 Experimental Setup
Dataset and Cohort Selection
We conduct our real data analysis using the Medical Information Mart for Intensive Care (MIMIC)-IV database [38, 39]. The MIMIC-IV dataset contains comprehensive information on over 40,000 patients admitted to intensive care units (ICUs) at the Beth Israel Deaconess Medical Center (BIDMC) between 2008 and 2019. This dataset includes a wide range of clinical data such as demographics, vital signs, laboratory results, procedures, diagnoses, and medications.
To select our cohort of patients with Acute Kidney Injury (AKI), we identified ICU stays where patients had ICD (International Classification of Diseases) codes linked to acute kidney failure, ensuring that the accompanying text explanations contained the keyword “acute kidney injury.” We excluded patients who, during their ICU stay, were labeled with burns, had end-stage renal disease, or renal dialysis. Additionally, patients with an estimated glomerular filtration rate (eGFR) of less than 15 mL/min/1.73 m2 were excluded to avoid including those with end-stage renal disease [40].
For data quality assurance, we removed ICU stays with fewer than 15 recorded measurements within the first 21 days after ICU admission. Furthermore, we excluded ICU stays that had two or fewer recorded measurements in each of the following intervals since ICU admission: 0–7 days, 7–14 days, and 14–21 days. The key feature used in our analysis was serum creatinine levels, a laboratory measurement that is the primary determinant for diagnosing and monitoring AKI [41].
Data Preprocessing and Model Selection
We perform unsupervised learning on the data. Specifically, we transform the original longitudinal data as detailed in Section 2, and used the estimated coefficients of the spline basis functions to represent each ICU stay. We applied $k$-means clustering [42] and $k$-medoids clustering [43] on these representations. The quality of the resulting clusters was evaluated by computing the average Silhouette coefficient [34].
For clinical evaluation, we labeled each ICU stay based on the AKI severity stages defined below [41, 44]:
  1. 1.
    Stage 1: Serum creatinine (SCr) level rises to 1.5×1.5$\times$1.5 × to 1.9×1.9$\times$1.9 × the baseline value within seven days since time 0.
  2. 2.
    Stage 2: SCr level rises to 2.0×2.0$\times$2.0 × to 2.9×2.9$\times$2.9 × the baseline value within seven days since time 0.
  3. 3.
[…]
    1. (a)
[…]
    2. (b)
      Subtype 2: SCr level rises to 3×3$\times$3 × or greater than the baseline value within seven days since time 0 and does not meet the criteria for Subtype 1.
The baseline SCr value was computed using the Modification of Diet in Renal Disease (MDRD) equation [45].
[…]
4.2 Unsupervised Learning – Phenotyping
We first evaluate the impact of timeline registration on clustering AKI ICU stays using $k$-medoids and $k$-means. The average Silhouette coefficients for different number of clusters $K\in[2,8]$ are presented in Table 2. Notably, when $K=2$ and $K=3$, the clustering results yield significantly higher average Silhouette coefficients, indicating more distinct clustering results. Additionally, we observe that both $k$-medoids and $k$-means show substantial increases in the average Silhouette coefficients after applying the proposed timeline registration method. Given the heterogeneity of AKI, we expect the ICU stays to cluster according to different subtypes of the condition. The numerical results demonstrate the effectiveness of our timeline registration method in aligning the timelines of patients within the same subtype.
$k$-medoids $k$-means
Observed Data Registered Data Registered Data Registered Data
K=2 0.431 0.483 0.584 0.628
K=3 0.226 0.260 0.523 0.545
K=4 0.249 0.269 0.325 0.332
K=5 0.208 0.280 0.350 0.380
K=6 0.208 0.271 0.352 0.395
K=7 0.111 0.112 0.346 0.402
K=8 0.108 0.159 0.284 0.220
Table 2: Comparison of average Silhouette coefficients for original and adjusted data after applying the proposed timeline registration method, using $k$-medoids and $k$-means clustering.
In computational phenotyping research, the number of subtypes $K$ is often unknown and must be determined. Therefore, we consider a scenario in which the number of clusters falls within a pre-specified range, and we apply the proposed timeline registration method to determine this number in a data-driven manner. In Table 3, we present results for two scenarios: one requiring a smaller number of clusters and the other requiring a larger number, as potentially suggested by auxiliary information. In both cases, we observe that data after timeline registration show increased average Silhouette coefficients, indicating that the timeline registration method provides consistent improvement in downstream clustering performance.
$k$-medoids $k$-means
Observed Data Registered Data Observed Data Registered Data
$K\in[2,4]$ 0.431 0.483 0.584 0.628
$K\in[5,8]$ 0.208 0.283 0.352 0.431
Table 3: Comparison of average Silhouette coefficients for original and adjusted data after applying the proposed timeline registration method, using $k$-medoids and $k$-means clustering. Number of clusters $K$ is either in the range of $[2,4]$ or $[5,8]$.
4.3 Clinical Evaluation
Next, we incorporate clinical information, specifically the AKI severity stages, into the evaluation of timeline registration performance. Recall that the definition of AKI severity stages, as outlined in Section 4.1, relies on the starting time point. This implies that the AKI severity stage of an individual subject may change after applying timeline registration. Therefore, for both the observed data (without timeline registration) and the data after applying the proposed timeline registration procedure, we label each subject as Stage 1 (S1), Stage 2 (S2), Stage 3 Subtype 1 (S3-1), or Stage 3 Subtype 2 (S3-2). We then group both the observed and registered datasets into five different configurations, as described in the first column of Table 4, and evaluate the average Silhouette coefficient for each configuration. The results for these five different AKI sub-category divisions are presented in Table 4. Consistently, the data processed through timeline registration demonstrate an increased average Silhouette coefficient, indicating improved data aggregation within each AKI sub-category.
Observed Data Registered Data
4-Groups (S1, S2, S3-1, S3-2) -0.049 -0.041
3-Groups (S1, S2, (S3-1, S3-2)) -0.057 -0.047
3-Groups ((S1, S2), S3-1, S3-2) 0.286 0.327
2-Groups ((S1, S2), (S3-1, S3-2)) 0.321 0.341
2-Groups ((S1, S2, S3-1), S3-2 ) 0.527 0.627
Table 4: Comparison of observed and recovered data after registration in the clinical setting using new criteria for defining creatinine baseline level.
4.4 Supervised Learning – Severe AKI Prediction
Observed Data Registered Data ($k$-medoids) Registered Data ($k$-means)
AUROC 0.8050 [0.7836, 0.8272] 0.8504 [0.8316, 0.8688] 0.8456 [0.8242, 0.8658]
Accuracy 0.8164 [0.8015, 0.8309] 0.8343 [0.8207, 0.8487] 0.8278 [0.8138, 0.8419]
AUPRC 0.8023 [0.7812, 0.8237] 0.8379 [0.8199, 0.8572] 0.8289 [0.8081, 0.8499]
From both tables, we observe that using data after registration results in significant increases in the AUROC (area under the receiver operating characteristic curve) and AUPRC (area under the precision-recall curve) values. Accuracy values also slightly increase. This comparison indicates that timeline registration is influential in enhancing downstream clinical analysis and can be considered an essential preprocessing step.
Observed Data Registered Data ($k$-medoids) Registered Data ($k$-means)
AUROC 0.8276 [0.8062, 0.8489] 0.8787 [0.8621, 0.8957] 0.8776 [0.8587, 0.8972]
Precision 0.8138 [0.7894, 0.8374] 0.8318 [0.8116, 0.8548] 0.8251 [0.8035, 0.8449]
Recall 0.7326 [0.7104, 0.7543] 0.7684 [0.7451, 0.7903] 0.7581 [0.7372, 0.7807]
Accuracy 0.8273 [0.8125, 0.8412] 0.8463 [0.8323, 0.8604] 0.8403 [0.8261, 0.8546]
AUPRC 0.8256 [0.8046, 0.8460] 0.8680 [0.8509, 0.8861] 0.8621 [0.8433, 0.8806]
5 Discussions and Conclusions
In this paper, we demonstrate the necessity of performing timeline registration for longitudinal EHR data, aligning patient records to reflect the intrinsic disease progression trends. To evaluate the feasibility of the proposed registration algorithm for recovering ground truth longitudinal data, we performed simulations with six different scenarios. The results indicate that the proposed method can effectively recover observed data to ground truth with low time complexity. To demonstrate the impact of registration on downstream clinical data analysis, we perform clustering with clinical evaluation and severe AKI prediction and compare their results using data with or without timeline registration. The results suggest that registration can effectively improve the performance of downstream clinical data analysis. We also conclude that the registration algorithm is of high stability, reflected by the minimal 95% confidence intervals given 500 times repetition of experiments. Additionally, our proposed registration algorithm has great generalizability in that its clustering step can be substituted by any desired clustering method. Registration on continuous space and evaluation of registration on other critical medical events will be investigated in our future work. EHR timeline registration is an emerging field and our work provides insight for improving clinical analysis using EHR data.

### v2

@@ A.3 Results
In summary, our algorithm effectively recovered the temporal positions from the simulated data across all eight scenarios. The evaluation metrics show similar values regardless of the complexity of the scenario, indicating the robustness of our algorithm. The total runtime per execution of the algorithm on the simulated dataset does not exceed 2 minutes, demonstrating its low computational complexity and potential for practical application on large-scale EHR datasets. We further compared our approach with dynamic time warping (DTW) Muller2007 under heterogeneous progression speeds. The results, presented in Section B of the Supplementary Materials, show the applicability of our method in a broader range of settings.
A.4 Sensitivity Analyses
Next, we assess the robustness of the proposed algorithm to common challenges in real-world EHR data, including outliers, missingness, and measurement noise. Here, sparsity and irregularity were modeled by randomly sampling observation time points from a uniform distribution over the interval $[1,21]$, with non-selected times considered missing. Noise was introduced using a random error term, $\epsilon_{ij}$, added to each simulated trajectory. We evaluate: (1) robustness to outliers by modifying the original simulation (Scenario 3, denoted SA0) to randomly select and double the values of 100, 200, and 300 observations, thereby introducing increasing levels of contamination (Figure 12); (2) robustness to missing data by randomly deleting 10%, 15%, and 20% of the data points at the generation stage (Figure 13); and (3) robustness to noise by increasing the variance of the additive Gaussian error term $\epsilon_{ij}$ by 15%, 30%, and 45%, respectively (Figure 14). Across all settings, the algorithm demonstrated stable performance.
Moreover, we conducted systematic sensitivity analyses to examine the impact of three key parameters, $M$, $\alpha$, and $\tau$, on the performance of the proposed algorithm. All analyses were performed under the simulation setting of Scenario 4 described in Section A in the supplementary materials.
Parameter $M$ represents the maximum number of candidate clusters considered after dimensionality reduction. Smaller values of $M$ restrict the search space, while larger values (even beyond the true number of subtypes) preserve performance but increase computational cost. When users have prior biomedical knowledge about the number of disease subtypes, we recommend choosing $M$ slightly above that expectation. Sensitivity analysis over values of $M$ ranging from 2 to 10 demonstrated that performance remains stable and optimal when $M$ is equal to or greater than the true number of subtypes (four), supporting our recommendation (Figure 15).
Parameter $\alpha$ controls the robustness of average trajectory estimates within each cluster by masking potential outliers during timeline alignment. While selecting $\alpha$ is challenging without prior knowledge of contamination, recent theoretical results under Huber’s contamination model suggest that fully adaptive methods may incur substantial statistical cost luo2024adaptive . Practically, $\alpha$ should not be set too high, as it may exclude too many data points and distort the representative trajectory. Sensitivity analysis across the range $[0.80,0.99]$ confirmed that the algorithm’s performance is robust to this parameter (Figure 16).
Parameter $\tau$ serves as a threshold for accepting the initial clustering, with higher values implying stronger confidence and fewer required iterations. In practical applications, users may refer to previously published recommendations, especially when using common public datasets. Additionally, general guidelines from the literature dalmaijer2022statistical suggest interpreting silhouette coefficients greater than $0.5$ as evidence for clustering and coefficients exceeding $0.7$ as strong evidence. However, silhouette coefficients exceeding $0.7$ are rarely observed in biomedical applications banerjee2023identifying ; manzini2022longitudinal ; coefficients around $0.5$ typically indicate sufficient quality. We conducted a sensitivity analysis over $\tau\in[0.45,0.55]$ and observed stable algorithmic performance across this range (Figure 17).
Figure 12: Validation of robustness to outliers. SA0 represents the original Scenario 3; the other settings were generated from SA0 by randomly selecting 100, 200, and 300 observations, respectively, multiplying them by 2, and thereby introducing progressively higher levels of outlier contamination. In each panel, the shaded band denotes the interquartile range (25th–75th percentiles) of the performance metric, and the central dot marks the median.
Figure 13: Validation of robustness to missing data. SA0 denotes the original Scenario 3; the other settings were created from SA0 during data generation by randomly deleting 10%, 15%, and 20% of the generated data points, thus introducing increasing degrees of missingness. In each panel, the shaded band denotes the interquartile range (25th–75th percentiles) of the performance metric, and the central dot marks the median.
Figure 14: Validation of robustness to noise. SA0 corresponds to the original Scenario 3; the other settings were generated from SA0 by increasing the variance of the Gaussian error term $\epsilon_{ij}$ by 15%, 30%, and 45%, respectively, thereby introducing higher noise levels. In each panel, the shaded band denotes the interquartile range (25th–75th percentiles) of the performance metric, and the central dot marks the median.
Figure 15: Sensitivity analysis on parameter $M$. The analysis is conducted under the simulation setting of Scenario 4; in each panel, the shaded band denotes the interquartile range (25th–75th percentiles) of the performance metric, and the central dot marks the median.
Figure 16: Sensitivity analysis on parameter $\alpha$. The analysis is conducted under the simulation setting of Scenario 4; in each panel, the shaded band denotes the interquartile range (25th–75th percentiles) of the performance metric, and the central dot marks the median.
Figure 17: Sensitivity analysis on parameter $\tau$. The analysis is conducted under the simulation setting of Scenario 4; in each panel, the shaded band denotes the interquartile range (25th–75th percentiles) of the performance metric, and the central dot marks the median.
A.5 Ablation Studies
1 66.8% [62.5%, 74.8%] 78.8% [72.8%, 87.4%] 0.62 [0.42, 0.79] 2.18 [0.92, 5.69]
2 64.6% [62.4%, 68.3%] 87.4% [77.2%, 91.6%] 0.54 [0.47, 0.73] 0.63 [0.50, 1.03]
3 67.5% [63.9%, 72.0%] 77.4% [72.5%, 85.4%] 0.66 [0.50, 0.79] 1.86 [0.98, 4.49]
4 67.1% [65.2%, 69.0%] 87.3% [85.3%, 89.6%] 0.52 [0.48, 0.57] 0.63 [0.50, 1.03]
5 68.1% [65.8%, 70.6%] 82.4% [75.8%, 88.2%] 0.63 [0.51, 0.77] 0.67 [0.51, 1.07]
6 67.1% [63.5%, 71.7%] 77.5% [73.6%, 84.6%] 0.66 [0.52, 0.76] 2.05 [1.01, 4.45]
7 61.6% [59.5%, 64.3%] 82.6% [79.9%, 87.3%] 0.69 [0.57, 0.76] 0.60 [0.53, 0.70]
8 66.3% [62.8%, 72.1%] 78.3% [73.1%, 85.0%] 0.68 [0.53, 0.77] 2.37 [0.97, 5.55]
Table 9: Ablation Studies: We present the average value and 95% confidence interval for the rate of exact recovery, rate of recovery error, mean absolute value, and runtime using the smoothing spline method.
1 55.7% [52.7%, 58.6%] 74.8% [71.6%, 79.2%] 0.88 [0.78, 0.96] 1.09 [0.67, 1.57]
2 49.9% [46.4%, 55.9%] 72.8% [69.4%, 77.6%] 0.97 [0.84, 1.06] 1.16 [0.70, 1.73]
3 53.4% [50.4%, 56.6%] 73.7% [70.3%, 78.3%] 0.93 [0.81, 1.02] 1.09 [0.69, 1.57]
4 50.2% [47.3%, 53.0%] 72.1% [68.1%, 76.2%] 0.98 [0.88, 1.09] 1.10 [0.70, 1.56]
5 53.3% [50.8%, 56.1%] 72.5% [68.9%, 76.8%] 0.95 [0.84, 1.04] 0.88 [0.77, 1.01]
6 54.4% [51.9%, 57.0%] 73.5% [70.1%, 77.8%] 0.92 [0.82, 1.01] 0.87 [0.76, 1.01]
7 46.1% [40.3%, 50.1%] 70.5% [66.3%, 73.8%] 1.06 [0.97, 1.18] 0.89 [0.79, 1.06]
8 53.2% [50.6%, 56.0%] 73.3% [69.8%, 77.8%] 0.94 [0.83, 1.02] 0.87 [0.76, 1.05]
Table 10: Ablation Studies: We present the average value and 95% confidence interval for the rate of exact recovery, rate of recovery error, mean absolute value, and runtime using the RKHS (Reproducing Kernel Hilbert Space) method.
1 66.0% [62.3%, 86.6%] 82.0% [76.9%, 99.0%] 0.58 [0.15, 0.69] 1.00 [0.52, 1.63]
2 72.8% [69.0%, 77.0%] 91.9% [88.3%, 95.0%] 0.41 [0.32, 0.52] 0.96 [0.52, 1.17]
3 73.3% [15.2%, 78.7%] 88.2% [61.9%, 91.1%] 0.45 [0.36, 1.46] 0.94 [0.52, 1.17]
4 73.5% [29.7%, 80.2%] 93.4% [80.0%, 96.9%] 0.39 [0.31, 1.05] 0.95 [0.52, 1.17]
5 70.4% [19.8%, 77.9%] 86.6% [62.4%, 92.6%] 0.58 [0.43, 1.34] 1.25 [0.52, 4.00]
6 77.2% [33.0%, 81.3%] 93.8% [85.3%, 95.6%] 0.38 [0.31, 1.00] 0.94 [0.52, 1.16]
7 70.6% [67.5%, 74.4%] 91.7% [90.5%, 93.8%] 0.47 [0.39, 0.53] 0.95 [0.51, 1.17]
8 69.9% [17.1%, 75.1%] 85.5% [48.0%, 88.9%] 0.54 [0.43, 1.87] 0.94 [0.51, 1.18]
Table 11: Ablation Studies: We present the average value and 95% confidence interval for the rate of exact recovery, rate of recovery error, mean absolute value, and runtime using the k-means method.
To assess the dimension reduction component, we fixed the clustering method (using $k$-medoids) and compared our primary cubic B-spline dimension reduction approach (Table 8) against two alternative dimension reduction methods: smoothing splines (Table 9) and an approach based on Reproducing Kernel Hilbert Space (RKHS, Table 10). All evaluated dimension reduction methods effectively recovered the true trajectories; the cubic B-spline approach consistently achieved the best overall performance, followed by the smoothing spline, and finally the RKHS-based method. These comparisons provide empirical justification for selecting cubic B-splines as our primary dimension reduction method.
To evaluate the clustering component separately, we fixed the dimension reduction method (cubic B-spline) and compared the performance of two clustering algorithms: $k$-medoids clustering (Table 8) and $k$-means clustering (Table 11). Both clustering methods yielded sufficiently good results, though $k$-medoids generally outperformed $k$-means. Further comparisons between these clustering methods were conducted in the real-data analyses, and the results are summarized in Tables 1–5 of the main text. Both clustering algorithms were effective; but $k$-medoids clustering exhibited greater stability in unsupervised tasks (Sections 4.2) and delivered superior performance in clinical evaluation and supervised prediction tasks (Sections 4.3 and 4.4). The greater robustness of the $k$-medoids algorithm likely contributes to this observed advantage, providing practical guidance for clustering algorithm selection in practice.
Appendix B Comparison with DTW Results
In this section, we compare the performance of our proposed registration method with the conventional dynamic time warping (DTW). Using the simulation setup described in Section A.1, we instantiate a scenario in which subjects exhibit heterogeneous disease progression speeds across subgroups. We consider two trajectory families: (i) a continuously heterogeneous family representing an imperfect subtyping setting, where some individuals are fast progressors (Group 2), some are slow progressors (Group 3), and the remaining display a continuum of intermediate speeds (Group 4); and (ii) a shape-heterogeneous family (Group 1), which shares a similar overall trend but differs in trajectory shape, again yielding imperfect subtyping relative to Groups 2–4.
The data generation setup is as follows. Recall that $s_{i}^{(\cdot)}t_{ij}$ denotes the observed time, and $s_{i}^{(\cdot)}$ represents the subject-specific progression speed factor:
Group 1 (300 subjects): $\quad Y_{i}^{(1)}(t_{ij})=7+1.5\,t_{ij}+4\sin\!\big(0.7(t_{ij}+4)\big)+\epsilon_{ij},$
Group 2 (250 subjects): $\quad Y_{i}^{(2a)}\!\big(s_{i}^{(2)}t_{ij}\big)=2\,t_{ij}+3\sin(t_{ij}+4)+\epsilon_{ij},$
Group 3 (250 subjects): $\quad Y_{i}^{(2b)}\!\big(s_{i}^{(3)}t_{ij}\big)=2\,t_{ij}+3\sin(t_{ij}+4)+\epsilon_{ij},$
Group 4 (200 subjects): $\quad Y_{i}^{(2c)}\!\big(s_{i}^{(4)}t_{ij}\big)=2\,t_{ij}+3\sin(t_{ij}+4)+\epsilon_{ij}.$
where $s_{i}^{(2)}=1$ for Group 2, $s_{i}^{(3)}=0.7$ for Group 3, and $s_{i}^{(4)}\sim\mathrm{Uniform}(0.7,1)$.
We evaluate three choices of the number of clusters, $K_{ture}\in\{2,3,4\}$, with the ground-truth subtypes defined as:
  • $K_{ture}=2$: Group 1 versus Groups 2–4, corresponding to two major subtypes with different shapes.
  • $K_{ture}=3$: Group 1 versus a relatively slow-progressor subtype versus a relatively fast-progressor subtype; the latter two constitute imperfect subtypes.
  • $K_{ture}=4$: Group 1, Group 2, Group 3, and Group 4, where Group 4 serves as a transitional subtype between Groups 2 and 3, again reflecting an imperfect subtyping configuration.
We then apply DTW and our registration method for downstream clustering and assess their ability to recover the true subtypes.
To evaluate clustering performance, we conducted 500 simulation replicates and computed three standard metrics: Adjusted Rand Index (ARI), Adjusted Mutual Information (AMI), and Clustering Accuracy (ACC). To ensure comparability across methods, we used a common protocol: (i) all clustering tasks were performed via $k$-medoids on the respective distance matrices; (ii) similar to the filtering in Section 4.1, we excluded individuals with no observations in either of the time windows $[0,8]$ or $(8,13]$ prior to computing the clustering metrics. These metrics respectively measure pairwise agreement, information overlap, and label correspondence between the true and predicted clusters. The average values and 95% confidence intervals for all three metrics over 500 runs across all replicates are summarized in Table 12.
$K_{ture}$ Metric DTW Our Method
2 ACC 0.697 [0.613, 0.748] 0.635 [0.615, 0.657]
AMI 0.078 [0.027, 0.214] 0.110 [0.083, 0.140]
ARI 0.112 [0.044, 0.198] 0.067 [0.045, 0.093]
3 ACC 0.439 [0.405, 0.487] 0.744 [0.722, 0.768]
AMI 0.098 [0.048, 0.167] 0.357 [0.317, 0.399]
ARI 0.079 [0.035, 0.146] 0.404 [0.363, 0.450]
4 ACC 0.431 [0.313, 0.507] 0.588 [0.563, 0.616]
AMI 0.206 [0.060, 0.348] 0.327 [0.290, 0.360]
ARI 0.170 [0.028, 0.291] 0.321 [0.293, 0.352]
Table 12: Comparison of DTW and our method across different $K_{ture}$ values (average value [95% CI])
Table 12 shows that when the number of clusters is small ($K_{ture}=2$), the two methods perform comparably. However, as the number of clusters increases ($K_{ture}=3$ and $K_{ture}=4$), our proposed registration method substantially outperforms DTW, achieving higher AMI and ARI scores and more stable clustering accuracy across replicates.
Appendix C Computational and Space Complexities
In this section, we analyze the computational complexity of the proposed framework by examining each step separately.
In Step 1 (Algorithm 1), each subject’s longitudinal trajectory is transformed into a set of basis coefficients via ridge regression for each candidate time shift in $\Lambda$. Incorporating $p$ basis functions, the ridge regression for each shift incurs a computational cost of $\mathcal{O}(T_{i}p^{2}+p^{3})$, where $T_{i}$ is the number of observations for subject $i$. Aggregating across all subjects and all shifts, the total complexity for Step 1 is approximately $\mathcal{O}(|\Lambda|(Tp^{2}+Np^{3}))$, where $T=\sum_{i=1}^{N}T_{i}$ denotes the total number of observations.
In Step 2 (Algorithm 2), an iterative clustering procedure with dynamic time shift updates is performed. In each outer iteration, the algorithm evaluates $M$ candidate clusterings using $k$-means, $k$-medoids, or a similar algorithm, and computes the silhouette score for each $k$. Assuming the clustering procedure converges in $I_{c}$ inner iterations, this step requires $\mathcal{O}(MNkpI_{c}+MN^{2}p)$ per outer iteration, where $p$ is the feature dimension. Additional operations such as centroid computation, trimmed averaging, and time shift updates for each subject together contribute $\mathcal{O}(Np+N|\Lambda|p)$ per iteration. With $I$ outer iterations, the total complexity for Step 2 is $\mathcal{O}(I(MNkpI_{c}+MN^{2}p+Np+N|\Lambda|p))$.
In Step 3 (Algorithm 3), unselected points are reassigned based on an exhaustive search over candidate shifts and clusters. For each unselected subject, this search requires evaluating $\mathcal{O}(|\Lambda|K^{(t_{\text{final}})}p)$ distances, where $K^{(t_{\text{final}})}$ is the number of clusters at termination. Assuming a constant fraction of subjects require reassignment, the total complexity for Step 3 is $\mathcal{O}(N|\Lambda|K^{(t_{\text{final}})}p)$.
Summing across all three steps, the dominant terms are $\mathcal{O}(N|\Lambda|Tp^{2})$ for Step 1 (since the number of basis functions $p$ is typically small and treated as a constant, e.g., around 10), $\mathcal{O}(IMN^{2}p)$ for Step 2 (since the number of subjects $N$ dominates the number of clusters, clustering iterations, and candidate shifts $|\Lambda|$), and $\mathcal{O}(N|\Lambda|Kp)$ for Step 3. Therefore, the overall computational complexity of the framework is $\mathcal{O}(N|\Lambda|Tp^{2}+IMN^{2}p)$, as the contribution from Step 3 is relatively minor when $T$ is much larger than $K$. Here, $N$ is the number of subjects, $T$ is the total number of observations, $p$ is the number of basis functions, $|\Lambda|$ is the number of candidate time shifts, $M$ is the maximum number of clusters considered, $K$ is the number of clusters selected at convergence, and $I$ is the number of outer iterations. The computational cost is dominated by the data transformation step when the number of time shifts or basis functions is large, and by the clustering step when the number of subjects is large.
On the other hand, for a given number of clusters $M$, the registration algorithm proposed in jiang2023timeline incurs a computational cost of $\mathcal{O}(INMT_{\text{total}}|\Lambda|)$, where $T_{\text{total}}$ denotes the total number of observations across the entire time grid. In the context of EHR data, this can be as large as $\left|\{t_{ik}\mid i=1,\ldots,N;\ k=1,\ldots,T_{i}\}\right|$. Furthermore, if we aim to determine the optimal number of clusters by using the silhouette index in practice, this introduces an additional computational cost of $\mathcal{O}(MN^{2}T_{\text{total}})$. In summary, the proposed algorithm yields a substantially lower computational cost compared to the method introduced in jiang2023timeline .
The space complexity of the proposed framework is primarily driven by the storage of transformed features and intermediate clustering outputs. In Step 1 (Algorithm 1), for each subject $i$ and each candidate time shift $l\in\Lambda$, the algorithm computes and stores a vector of $p$ spline coefficients. This leads to a three-dimensional tensor $\mathbf{\Omega}$ of size $N\times|\Lambda|\times p$, contributing a space complexity of $\mathcal{O}(N|\Lambda|p)$. In Step 2, only the coefficients corresponding to the currently selected shift $\delta_{i}^{(t)}$ for each subject are used in clustering, so no additional large data structures are required beyond temporary arrays of size $\mathcal{O}(Np)$ for centroid and distance calculations. Step 3 similarly reuses the stored coefficient tensor and does not introduce significant new memory demands. Therefore, the overall space complexity of the algorithm is $\mathcal{O}(N|\Lambda|p)$, where linear in the number of subjects, candidate shifts, and basis coefficients. This cost can be significantly lower than the space complexity of the method in jiang2023timeline , which is $\mathcal{O}(NT_{\text{total}})$.
In summary, our analysis of computational and space costs demonstrates the feasibility of applying our method to large-scale EHR datasets.